OpenStudy (calculusxy):

Evaluate the function f(x) = 1/(x-4) when f(1/x)

OpenStudy (calculusxy):

@phi

jhonyy9 (jhonyy9):

substitute the 1/x in place of x inside equation - try it and see what will get ?

jhonyy9 (jhonyy9):

f(1/x) = ?

OpenStudy (calculusxy):

so we will have \(\large \frac{1}{\frac{1}{x} -4}\)

jhonyy9 (jhonyy9):

yes and continue it

OpenStudy (calculusxy):

i don't know how to evaluate it

jhonyy9 (jhonyy9):

first 1/x -4 = ?

OpenStudy (calculusxy):

\[\frac{ 1 }{ x } - 4\] i don't know what I am supposed to do. please help me in step-by-step

jhonyy9 (jhonyy9):

how you assum a fraction with an integer ?

jhonyy9 (jhonyy9):

common denominator ?

jhonyy9 (jhonyy9):

1 4 --- - ----- = ? x 1

OpenStudy (calculusxy):

the common denominator would be x?

jhonyy9 (jhonyy9):

yes

OpenStudy (mathmale):

Your "when f(1/x)" is unclear. What are you trying to say here? You do not yet have an x-value to substitute into the first expression. Please ensure you've copied down all parts of this problem. Thx.

OpenStudy (calculusxy):

1/x - 4x/x = (1 - 4x)/x

OpenStudy (calculusxy):

@mathmale the input value is (1/x)

OpenStudy (mathmale):

OK. Leave out the "f" in the second expression. You could write, "Evaluate the function f(x) = 1/(x-4) when the input to f(x) is 1/x."

jhonyy9 (jhonyy9):

@calculusxy yes is right but not forget that there is 1/---

OpenStudy (mathmale):

In that case you'd want to evaluate f(1/x); the word "when" should not be there.

OpenStudy (calculusxy):

@jhonyy9 so what do i do with that numerator?

jhonyy9 (jhonyy9):

check what you have got for denominator than rewrite it like how have got firstly using 1/ --- and hope you know that 1 d ------ = 1*----- = ? a+b a+b ----- d hope this is understandably

OpenStudy (mathmale):

Evaluate the function f(x) = 1/(x-4) .... Eliminate the "x" from this expression, replacing it with ( ): f( ) = 1 / ( ( ) -4). Finally, write the input "1/x" inside the ( ) and simplify. This is the same as jhonny9 suggested 'way back: "jhonyy9 substitute the 1/x in place of x inside equation - try it and see what will get ?"

OpenStudy (mathmale):

"Evaluate the function f(x) = 1/(x-4) when f(1/x)" would be clear if re-written as Evaluate f(1/x) when the function f(x) is defined as f(x) = 1/(x-4)."

OpenStudy (calculusxy):

@jhonyy9 i have the expression \(\large \frac{1}{\frac{1 - 4x}{x} - 4}\). is the denomintor correct?

OpenStudy (calculusxy):

wait. get rid of the -4

jhonyy9 (jhonyy9):

no - without -4 because you have added to 1/x -4 and have got

OpenStudy (calculusxy):

\(\large \frac{1 - 4x}{x}\)

jhonyy9 (jhonyy9):

yes so

jhonyy9 (jhonyy9):

yes and continue it

OpenStudy (phi):

fyi, once you wrote \[ \large \frac{1}{\frac{1}{x} -4} \] that is the answer. Anything after that is changing its form to make it easier to use (or to match it up with an answer choice)

OpenStudy (calculusxy):

@phi I actually got \(\large \frac{1}{\frac{1 - 4x}{x}}\)

OpenStudy (calculusxy):

do i have to do anything with the complex fraction?

jhonyy9 (jhonyy9):

so and you can simplifie it again

OpenStudy (calculusxy):

okay so it would be \[\frac{ 1 }{ 1 } \times \frac{x}{1- 4x} = \frac{x}{1-4x}\]

OpenStudy (phi):

generally people don't like fractions inside fractions, so I would multiply top and bottom by x to clear the denominator.

jhonyy9 (jhonyy9):

yes @calculusxy so is right sure

OpenStudy (calculusxy):

so my answer is \(\large \frac{x}{1 - 4x}\)

jhonyy9 (jhonyy9):

yes - right

OpenStudy (calculusxy):

thanks!

jhonyy9 (jhonyy9):

np was my pleasure good luck bye